# Predicting Multi-Agent Specialization via Task Parallelizability

### v1

@@ 2.2 Jensen-Shannon Divergence and Specialization
Let the set of high-level actions be denoted as $\mathcal{A}=\{a_{1},a_{2},...,a_{k}\}$ (Amato et al., 2019). Each agent follows a state-visitation distribution that describes how frequently they occupy different states under their policy (Schulman et al., 2015) given by:
@@
This distribution captures both what actions an agent tends to take and how often they visit states where those actions are available. We approximate specialization using the Jensen-Shannon divergence (JSD) over agents’ action distributions $P_{i}$ (Endres & Schindelin, 2003; Fuglede & Topsoe, 2004). Given $N$ agents with action distributions $P_{1},P_{2},...,P_{N}$, JSD is given by:
$\text{JSD}(P_{1},P_{2},...,P_{N})=\sum_{i=1}^{N}w_{i}D_{\mathrm{KL}}(P_{i}||M),$ (4)
where $M=\sum_{i=1}^{N}w_{i}P_{i}$ is the mixture distribution, $w_{i}=\frac{1}{N}$, and $D_{\mathrm{KL}}(P_{i}||M)$ is the Kullback-Leibler (KL) divergence between each agent’s action distribution $P_{i}$ and $M$. When $\text{JSD}(P_{1},P_{2},...,P_{N})=0.0$, all $N$ agents have performed the same proportions of actions, indicating that all agents have performed all subtasks equally (no specialization). When $\text{JSD}(P_{1},P_{2},...,P_{N})=1.0$, the action distributions are completely disjoint (i.e., each agent assigns probability exclusively to a unique set of actions with no overlap), meaning that the agents specialize in complementary subtasks. Intermediate values of JSD indicate partial specialization, in which the agents overlap in some subtasks but not others. We use Jensen-Shannon rather than Kullback-Leibler divergence because it is symmetric, i.e. $\text{JSD}(P_{1}||P_{2})=\text{JSD}(P_{2}||P_{1})$. The degree of specialization in each team should be the same regardless of which agent specializes in which subtasks.
2.3 Graphical Representations of Spatial Layouts
To model task execution in structured environments, we represent spatial layouts as graphs. This allows for efficient computation of movement paths, bottlenecks, and spatial constraints on multi-agent coordination. We define a layout graph $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of walkable positions and task-relevant objects and $\mathcal{E}$ is the set of valid movements between adjacent positions. This spatial representation applies broadly to spatio-temporal multi-agent environments, including gridworlds. $G$ defines the static structure of the environment and constrains the MDP, where the state space $\mathcal{S}$ includes agent positions $x\subseteq\mathcal{V}$ and task-relevant object states. The transition function $\mathcal{T}$ ensures that agents execute valid movements along edges in $\mathcal{E}$.
Definition 2.1.
@@ 2.3 Graphical Representations of Spatial Layouts
$V=\{u\mid u\in\{(i,j)\mid i\in[1,n],j\in[1,m]\}\setminus\text{obstacles}\}$
@@
2.4 Task Decomposition as a Directed Acyclic Graph
To analyze task parallelizability, we represent task execution as a directed acyclic graph (DAG), where nodes denote subtasks and edges represent the order in which subtasks must be completed (Solway et al., 2014).
Definition 2.2.
We define a task graph $T=\{\mathcal{B}_{r},\mathcal{E}_{T}\}$, where $\mathcal{B}_{r}$ (subtask nodes) represent intermediate steps such as workstations or resources, and $\mathcal{E}_{T}$ indicate ordering constraints between subtasks (e.g., an agent must retrieve an item before assembling it).
The number of subtasks $m$ depends on task complexity. Using the layout graph $G$, we determine the spatial locations and paths between subtasks. The fraction $f(i)$ of the total task spent on a given subtask $i$ is:
$f(i)=\frac{d(i)}{\sum_{j=1}^{m}d(j)},$ (5)
where $d(i)$ is the path length required to complete subtask $i$.
2.5 Spatial and Resource Bottlenecks
We introduce a framework to identify bottlenecks that constrain parallelizability in multi-agent tasks. This idea is inspired by sequential dependencies, resource contention, and communication overhead that prevent scalability in parallel processing (Almasi & Gottlieb, 1994; McCool et al., 2012; Zhuravlev et al., 2010). A classic example is memory bandwidth in GPUs, where increasing the number of threads does not lead to proportional speed-up due to limited shared access (Dublish et al., 2017). We extend this framework to multi-agent tasks by identifying two key types of bottlenecks: (1) spatial bottlenecks ($\mathcal{B}_{s}$) limit movement of multiple agents due to obstacles, narrow passages, and congestion, and (2) resource bottlenecks ($\mathcal{B}_{r}$) limit access to critical resources such as workstations or tools needed to complete subtasks (Mieczkowski et al., 2024). Together, these bottlenecks determine how effectively agents can execute tasks in parallel. They arise across numerous multi-agent domains, including warehouse operations (insufficient space, storage locations, and pallets (Živičnjak et al., 2022)), autonomous vehicles (lane reductions, traffic congestion (Zhang & Gao, 2020)), and smart factories (Wang et al., 2016).
@@ 2.5 Spatial and Resource Bottlenecks
We represent spatial and resource bottlenecks as two global sets over $G=(\mathcal{V},\mathcal{E})$: $\mathcal{B}_{s}\subseteq\mathcal{E}$, $\mathcal{B}_{r}\subseteq\mathcal{V}$, where $\mathcal{B}_{s}$ consists of edges in $\mathcal{E}$ that constrain movement and $\mathcal{B}_{r}$ consists of nodes in $\mathcal{V}$ with task-relevant objects (e.g., workstations). Each subtask $i$ is constrained by a subset of bottlenecks relevant to the path ($\mathcal{B}_{s}(i)$) and resources ($\mathcal{B}_{r}(i)$) required to complete it.
@@
To quantify spatial bottlenecks, we compute edge betweenness centrality, a graph-theoretic measure of congestion where high-centrality edges indicate critical bottlenecks in connectivity (Barrat et al., 2004).
Definition 2.3.
The betweenness centrality of an edge $e$ in graph $G=(\mathcal{V},\mathcal{E})$ is:
$C_{B}(e)=\sum_{s\neq t\in V}\frac{\phi(s,t|e)}{\phi(s,t)},$ (6)
where $\phi(s,t)$ is the number of shortest paths between nodes $s$ and $t$, and $\phi(s,t|e)$ is the number of those paths that pass through edge $e$.
@@
Since not all bottlenecks affect every subtask, we compute a per-subtask spatial bottleneck score by summing the centralities of all edges along the shortest path required for that subtask: $\mathcal{B}_{s}(i)=\sum_{e\in\text{path}(u,v)}C_{B}(e)$, where path$(u,v)$ is the shortest path between subtask-relevant nodes $u$ and $v$. A higher $\mathcal{B}_{s}(i)$ indicates stronger spatial constraints that may inhibit parallel agent movement.
@@
To quantify resource bottlenecks, we simply sum the number of task-relevant nodes that are required for a subtask $i$. For example, in a manufacturing system there may be only two welding stations; thus, only two agents may complete that subtask at the same time. Formally, for each subtask $i$, we define a resource bottleneck score $\mathcal{B}_{r}(i)$ as: $\mathcal{B}_{r}(i)=\sum_{v\in\mathcal{B}_{r(i)}}c(v)$, where $c(v)$ is the capacity of each task-relevant node (or the maximum agents that can use it at once).
2.6 Task Parallelizability Prediction
We now adapt Amdahl’s Law, a measure of the expected speed-up when dividing a task across multiple processors given the proportion that must be executed serially, to multi-agent tasks (Amdahl, 1967; Hill & Marty, 2008; Cassidy & Andreou, 2011). For the original form of Amdahl’s Law, see Appendix A.1. Mieczkowski et al. (2024) extended this idea to predict idleness during human collaborations given the current group size, workload, and resource constraints. We hypothesize that a similar formulation can be used to predict specialization in multi-agent tasks.
Let $f(i)$ represent the fraction of the total task time required for the $i^{th}$ of $m$ subtasks such that $\sum_{i=1}^{m}f(i)=1$. For each subtask $i$, the overall bottleneck capacity is defined as the minimum bottleneck that prevents agents executing it concurrently, or $\mathcal{B}_{i}=\min(\mathcal{B}_{s}(i),\mathcal{B}_{r}(i))$. The subtask speed-up factor $s_{i}(N,\mathcal{B}_{i})$ quantifies the efficiency gain from adding an $N$th agent to a subtask, given that $N-1$ agents are already working on it and $\mathcal{B}_{i}$ bottlenecks restrict parallel execution. Formally:
$s(N,\mathcal{B}_{i})=\frac{\text{min}(N,\mathcal{B}_{i})}{\text{min}(N-1,\mathcal{B}_{i})}$ (7)
Thus, if we assume that there is no cost to switching between subtasks, the overall expected speed-up achievable when an additional agent performs the entire task in parallel, which we call task parallelizability, is expressed as:
$S(N,\mathcal{B})=\frac{1}{\sum_{i=1}^{m}\frac{f(i)}{s(N,\mathcal{B}_{i})}}$ (8)
2.7 Task Parallelizability and Specialization
Our task parallelizability prediction estimates the efficiency gain if an additional agent independently completed all $m$ subtasks in parallel. We propose that this measure sets a fundamental upper bound on the benefits of specialization in multi-agent systems. If $S(N,\mathcal{B})$ is low due to resource or spatial bottlenecks, then the task would not be significantly sped up by multiple agents executing it in parallel. In this case, there would be little value to a team of generalist agents, so if we want to increase task efficiency it is better for agents to specialize and perform complementary subtasks. However, specialization itself can introduce new bottlenecks, particularly when agents must coordinate frequently or wait for specific subtasks to be completed before proceeding. In environments with high parallelizability $S(N,\mathcal{B})$, generalist agents who work independently may avoid these delays, leading to faster performance.
3 Application to Overcooked-AI
To investigate how task parallelizability influences specialization, we apply our theoretical model for predicting generalist efficiency to a specific multi-agent environment (Carroll et al., 2019). Overcooked-AI is a fully cooperative two-agent platform in which pairs of agents must deliver soups as quickly as possible while navigating obstacles, avoiding collisions, and coordinating subtasks (Wu et al., 2021). Each recipe requires a different number of subtasks ($m$) to prepare a soup, while resource bottlenecks such as the number of pots ($\mathcal{B}_{r}$) and spatial bottlenecks such as narrow passages ($\mathcal{B}_{s}$) constrain task parallelizability.
Layout Graphs. We represent each layout graph according to Definition 2.1. An example is depicted in Figure 2.
Figure 2: Examples of Overcooked layouts, layout graphs, and spatial bottleneck scores $\mathcal{B}_{s}(i)$ for subtask Onion $\to$ Pot. Layout A is an open room with low spatial bottlenecks, whereas B has a narrow passage leading to the pot, creating a greater bottleneck.
Task Decomposition. We decompose the overall task of “cooking a soup” using a directed acyclic graph (DAG) $T=(\mathcal{B}_{r},\mathcal{E}_{T})$ according to Definition 2.2, where $\mathcal{B}_{r}$ is the set of resources, or workstations (e.g., onion piles, pots, bowls), and $(m_{i}\to m_{j})\in\mathcal{E}_{T}$ is a directed edge representing a required subtask (such as moving an onion to the pot). An example can be found in Figure 5. The number of subtasks $m$ depends on the recipe. To make a one-onion soup, there are $m=4$ subtasks total: Onion $\to$ Pot, Cook in Pot, Bowl $\to$ Pot, and Pot $\to$ Serve. To make a two-onion or one-onion one-tomato soup, there are $m=5$ subtasks. To make a three-onion or two-onion one-tomato soup, there are $m=6$ subtasks. Using the layout graph $G$, we determine the locations and paths between workstations. The fraction $f(i)$ spent on subtask $i$ is its path length relative to the total path length of all subtasks following Equation 5.
Bottlenecks. We manipulate and measure bottlenecks in Overcoooked in two ways. We vary resource bottlenecks $\mathcal{B}_{r}$ by changing the number of workstations ($p$ pots, $o$ onions, $t$ tomatoes, $d$ bowls, and $s$ serving stations) available to agents. We vary spatial bottlenecks $\mathcal{B}_{s}$ by modifying the layout size and arrangement of obstacles (represented as counters that agents cannot cross). We quantify these spatial bottlenecks for each subtask $i$ using edge betweenness centrality as described in Definition 2.3.
Specialization. To measure specialization, we convert agents’ behaviors (move, interact, wait) from a policy roll-out of the best-performing seed to high-level actions: (i) picking up onions, (ii) placing onions in pot, (iii) picking up tomatoes, (iv) placing tomatoes in pot, (v) picking up bowls, (vi) moving soup from pot to bowl, and (vii) serving soup. Each agent $i\in\mathcal{N}$ has an action distribution $P_{i}:\mathcal{A}\to[0,1]$, where $P_{i}(a)$ represents the probability of agent $i$ taking action $a\in\mathcal{A}$ during a policy roll-out after training. For each agent $i$, we compute their high-level action distribution $P_{i}$ by normalizing the frequency of each action over the total actions they performed during the roll-out. We measure specialization as the Jensen-Shannon divergence between $|\mathcal{N}|=2$ agents, $\text{JSD}(P_{1}||P_{2})$.
Task Parallelizability Prediction. Our model for predicting task parallelizability measures the relative efficiency gained from two generalist agents making soups in parallel versus specializing in subtasks (e.g., fetching ingredients versus serving soup). Consider the task of making one-onion soup. This requires $m=4$ subtasks: Onion $\to$ Pot, Cook in Pot, Bowl $\to$ Pot, and Pot $\to$ Serve. Each subtask $i$ is associated with (i) a number of workstations $\mathcal{B}_{r}(i)\in\mathcal{B}_{r}$, (ii) a path length $d(i)$, and (iii) a measure of spatial bottlenecks $\mathcal{B}_{s}(i)$ between those workstations, where $\mathcal{B}_{s}(i)\geq 0.8$ indicates constrained areas. If there are multiple paths, we consider $d_{\min}(i)$. Subtask durations are weighted by path length (Equation 5), with Cook assumed to take 10 time-steps. The overall speed-up expected from two generalist agents making soups in parallel is thus given by $S(N,\mathcal{B})$ in Equation 8. We hypothesize an inverse relationship between $S(N,\mathcal{B})$ and $\text{JSD}(P_{1}||P_{2})$. This would suggest agents specialize more when task and environmental constraints inhibit them from executing the same subtasks simultaneously.
4 Experiment 1: Exploring Specialization in Large-Scale Overcooked Simulations
Figure 3: Experiment 1 Results. (A) Histogram of specialization (JSD) in the best-performing seeds for 3,200 unique Overcooked configurations reveals a bimodal distribution, with teams clustering as either fully generalist or specialist. (B) Scatter plot of predicted task parallelizability ($S$) vs. observed JSD reveals a moderate negative correlation, indicating less specialization when a task has greater parallelizability. (C) Variables influencing parallelizability show strong, significant effects on JSD: specialization decreases with more pots but increases with greater spatial bottlenecks and task complexity. Error bars represent standard error.
In Experiment 1, we conduct a large-scale analysis across a variety of environments to test whether parallelizability impacts the emergence of specialization.
4.1 Experiment Design
We generated a test suite of 3,200 unique Overcooked environments for exploratory analysis. We utilized 14 spatial configurations representing 2D gridworld kitchens, each with a unique arrangement of counters and floor areas, including the original layouts from Overcooked-AI and a new set of layouts inspired by real-world kitchens (Pejic et al., 2019). A complete set of layouts is provided in Figure 6. To create diverse parallelizability constraints, we systematically varied the number of workstations (onions, tomatoes, pots, bowls, and serving stations). A total of 16 combinations included one or two onion piles, bowls, pots, serving stations, and zero or one tomatoes. Workstation positions were randomized three times per layout. Agents were trained to make five recipes: (i) 1-onion (ii) 2-onion (iii) 3-onion (iv) 1-onion and 1-tomato and (v) 2-onion and 1-tomato soup. These variations produced a total of 3,200 unique environments. For each of these environments, we trained two independent agents to play Overcooked collaboratively via self-play using PureJaxRL and PPO with five random seeds (Schulman et al., 2017; Lu et al., 2022). Full details regarding training procedures can be found in Appendix A.3. We then selected the highest-performing seed based on the joint reward that agents achieved during the roll-out. If multiple seeds reached the same maximum reward, we averaged their JSD. Trials were excluded if agents failed to take high-level actions or achieved a roll-out reward of 0.
4.2 Results
Distribution of JSD. Results from Experiment 1 are depicted in Figure 3. Our simulations revealed two distinct behavioral modes. Specialists ($JSD\geq 0.5$) dominated $74.9\%$ of runs, with $48.9\%$ fully specialized ($JSD>0.9$). Generalists ($JSD<0.5$) appeared in $25.1\%$ of runs, including $22.3\%$ fully generalist ($JSD<0.1$). The remaining $28.8\%$ exhibited mixed specialist-generalist behaviors ($0.1\leq JSD\leq 0.9$).
Effects of Parallelizability on Specialization. For each of the variables presented in our predictive model, there was a significant effect in the expected direction on observed JSD. Full statistical details can be found in Appendix A.4. We found a significant effect of the number of pots, but not the total number of workstations; spatial bottlenecks; and recipe complexity (the number of ingredients required to make a soup) on emergent specialization.
Model Evaluation. We analyzed the relationship between predicted task parallelizability and observed specialization (JSD) across 14 Overcooked layouts. Full statistical details are in Appendix A.5, with Tables 1 and 2 summarizing correlations, JSD, and rewards. The average correlation $-0.486$ ($p=8.49\times 10^{-28}$) revealed a moderate negative relationship between generalist efficiency and JSD, supporting our hypothesis that greater task parallelizability predicts lower specialization. A logistic regression classifying low vs. high specialization ($JSD\geq 0.5$) achieved $74.2\%$ accuracy, with $0.93$ precision ($F1=0.81$) for high specialization and $0.46$ precision ($F1=0.58$) for low specialization.
Effects of State Space on Specialization We also observed a moderate correlation between grid size and average JSD ($r=0.383,p=5.57\times 10^{-17}$), even when controlling for resource and spatial bottlenecks (Figure 7). This suggests that state space size and workstation distances may act as confounding variables, influencing specialization independently of task-specific constraints. Larger state spaces increase exploration demands. In response, agents may adopt specialization as a locally optimal strategy to reduce policy complexity by focusing on specific paths and workstation interactions.
5 Experiment 2: Validating Predictive Model
Figure 4: Experiment 2 Results. (A) Six layouts (four depicted) with varying bottlenecks were designed on a $5\times 5$ grid, each with one or two pots, two onions, one bowl, and one serving station across four locations, yielding 48 layouts. Each layout was paired with one-, two-, or three-onion soup recipes for a total of 144 trials. (B) Histogram of specialization (JSD) across the best-performing seeds reveals a distinctly bimodal distribution, with most clustering around low (generalist) or high (specialist) JSD. (C) Scatter plot of predicted task parallelizability ($S$) vs. observed JSD averaged. A strong negative correlation confirms that greater potential speed-up from generalists corresponds to lower specialization.
In Experiment 1, we confirmed that the variables which should affect generalist strategies from speeding up task completion have a significant impact on specialization. In Experiment 2, we validate our results using a smaller, controlled experiment. We also show that we can use our findings to design environments which will induce or inhibit the emergence of specialized roles.
5.1 Experiment Design
We utilized the same methods, game dynamics, and training schemes as in Experiment 1 (described in Section 4.1). To control state space size while varying spatial bottlenecks, we constructed six $5\times 5$ grid layouts: (1) open space with surrounding counters, (2) single vertical divider, (3) double vertical divider and (4) horizontal divider. These spatial layouts are illustrated in Figure 4. We varied the number of pots (1 or 2) rather than workstations, since pots had the largest effect on specialization in Experiment 1, while controlling for relative distances between workstations across trials. Each layout featured four workstation arrangements, totaling 48 configurations. For each, agents were trained to make one-, two-, or three-onion soup with 10 random seeds for a total of 1400 runs. As in Experiment 1, we again selected the highest-performing seed in terms of the joint reward that agents achieved during the policy roll-out after averaging over the randomized locations of workstations. Four trials were excluded from Experiment 2.
5.2 Results
Results from Experiment 2 are depicted in Figure 4. By fixing the size of the grid, the distinction between generalist and specialist strategies became more pronounced, with a clear bimodal distribution emerging. There were 81 specialist teams with $JSD\geq 0.5$ ($57.86\%$) and 59 generalist teams ($42.14\%$). Notably, the proportion of generalists increased substantially compared to Experiment 1 once we removed the confound of altering layout size, providing further evidence that specialization is partially driven by the size of the state space. Additionally, our predictive model achieved even greater accuracy for predicting specialization. The Pearson correlation between predicted task parallelizability and JSD was $-0.667$ ($p=8.94\times 10^{-6}$), confirming the hypothesized inverse relationship between generalist efficiency and specialization. Statistical details can be found in Appendix A.6. To validate the model’s predictive power, we fit a class-balanced logistic regression to classify high ($JSD\geq 0.5$) or low specialization ($JSD<0.5$). The model achieved a test set accuracy of $90.90\%$, demonstrating strong performance with $0.88$ precision ($F1=0.93$) for high specialization and $1.00$ precision ($F1=0.86$) for low.
These results suggest that spatial bottlenecks and resource availability are key factors driving the emergence of specialization. Our findings indicate that our predictive model can be used to systematically design environments that will either promote or inhibit specialization. Rather than tailoring training algorithms to induce policy diversity, it is possible to do so based on environmental design alone. If agents are given ample space and workstations to learn the task, they will develop emergent generalist behaviors. Alternatively, if the environment is constrained spatially or in terms of resources, agents will develop specialist roles.
6 Experiment 3: Investigating State Space Size and Specialization
In Experiment 3, we investigate the relationship between state space size and specialization in greater detail. Our findings show that as the state space expands, even if the size of the grid that the agents can explore is held constant, agents consistently converge on more specialized strategies, indicating that specialization is a less computationally costly alternative when learning is made more difficult. These findings highlight a challenge in MARL: as state spaces grow, agents struggle to converge on optimal generalist policies. Full details, including the experimental design, results and statistical analysis are provided in Appendix A.7 and A.8.
7 Related Work
The study of specialization in multi-agent systems spans multiple disciplines. Division of labor leading to functional specialization has been widely studied in biological and social domains, including cells (Bell & Mooers, 1997; Rueffler et al., 2012), insect colonies (Ratnieks & Anderson, 1999; Fjerdingstad & Crozier, 2006), and human societies (Kitcher, 1990; Kreider et al., 2022; Goldstone et al., 2024; Vélez et al., 2024). Inspired by these insights, research in MARL and robotics has sought to design algorithms that actively encourage specialization by promoting policy diversity (Murciano et al., 1997; Padgham & Winikoff, 2002; Zhu & Zhou, 2008; Charakorn et al., 2020; Wang et al., 2020a, b; Li et al., 2021; Bettini et al., 2023, 2024) or by organizing agents into hierarchies (Zhang et al., 2010; Ahilan & Dayan, 2019). Similarly, new work on language-agent collaboration often assumes that specialization is necessary, prompting agents to adopt distinct roles before collaborating (Wang et al., 2023; Suzgun & Kalai, 2024; Swanson et al., 2024; Liu et al., 2024; Juang et al., 2024). However, much of this work assumes that specialization is always beneficial, without systematically identifying when and why generalist strategies may outperform specialists. A useful contrast comes from distributed systems, which are explicitly engineered rather than emergent. In these systems, heterogeneity and parallelism are carefully designed to optimize performance (Moncrieff et al., 1996; Almasi & Gottlieb, 1994; Pancake, 1996) and minimize contention (Park & Reveliotis, 2001). This structured approach provides a useful reference point for understanding when emergent specialization is not just incidental, but optimal.
While biological and social systems provide compelling examples, it remains an open question when specialization emerges as the optimal architecture. Prior work suggests that under conditions of high environmental variability or individual failure, flexible generalist strategies may outperform rigidly specialized roles (Staps & Tarnita, 2022; Ben-Oren et al., 2023). In our work, we propose a novel theoretical framework for predicting the utility of specialization based on task efficiency, considering environmental bottlenecks, resource constraints, and task complexity. Prior work in RL has shown that spatial bottlenecks such as doorways can shape subgoals (Menache et al., 2002), while research in human-robot collaboration and MARL has demonstrated that environmental design influences emergent behavior (Fontaine et al., 2021; McKee et al., 2022). Building on these insights, we formalize a predictive model for when generalist versus specialist strategies should emerge.
8 Discussion
Specialization, or the differentiation of agents’ policies and roles, is often assumed to improve collaborative performance. However, principles from distributed systems offer an alternative hypothesis. Specifically, generalist agents may outperform specialist ones when concurrent execution of the entire task would improve speed and efficiency. In this work, we outlined a theoretical framework for predicting generalist versus specialist efficiency in cooperative multi-agent tasks inspired by a variant of Amdahl’s Law. We then validated this model across three Overcooked experiments, finding highly significant effects of parallelizability factors on emergent specialization and high predictive accuracy. Together, these findings suggest that specialization is constrained by the environment and parallelizability of a task, making it possible to design layouts which either encourage or inhibit role emergence regardless of whether the training algorithm explicitly incentivizes diversity. More broadly, this framework motivates the use of distributed systems in analyzing trade-offs that arise when dividing labor.
This work has several limitations. First, our model does not account for switching behavior (generalists who switch roles rather than cooking independently) or cases where specialist and generalist policies are equally optimal. This limitation arises from an implicit assumption from Amdahl’s Law that subtasks are executed sequentially without overlap. Second, our model provides a theoretical upper bound on the benefits of generalists, but does not account for potential coordination overhead when increasing the number of agents, which could make the empirical relationship non-monotonic. Third, our model assumes agents are equally capable, yet specialization can also arise when agents have heterogeneous skills that make them more efficient at certain tasks (Moncrieff et al., 1996; Ginting et al., 2022). Despite these limitations, our predictive model is highly extensible. It generalizes from 2 to $n$ agents, providing explicit predictions for larger teams. Finally, our work suggests that specialization is a locally optimal solution that agents use when exploration is costly, i.e. in larger state spaces. Clarifying this trade-off can provide insight into how specialization may improve sample efficiency, as well as how to decrease coordination overhead between agents.
Acknowledgments
This work was supported in part by the Department of Defense (DoD) through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, the Office of Naval Research (ONR) MURI grant (N00014-22-1-2813), the EPSRC CDT in RAS (EP/L016834/1), and a grant from the Templeton World Charity Foundation.
Appendix A Appendix
A.1 Original Form of Amdahl’s Law
Let $f$ represent the proportion of a task that can be parallelized and accelerated by a factor of $s$, while $1-f$ denotes the portion that must be executed serially. The theoretical speedup $S$ achievable when distributing a task across multiple processors is given by:
$S=\frac{1}{(1-f)+\frac{f}{s}}$ (9)
A.2 Task Decomposition Graph
A.3 Training Procedures

### v2

@@ Appendix A State-Visitation and Action Distributions
In order to measure the differences between agents’ policies, we must quantify how different their distributions of states and actions are. Each agent follows a state-visitation distribution that describes how frequently they occupy different states under their policy [26] given by:
@@
This distribution captures both what actions an agent tends to take and how often they visit states where those actions are available.
Appendix B Original Form of Amdahl’s Law
Let $f$ represent the proportion of a task that can be parallelized and accelerated by a factor of $s$, while 1f1-f denotes the portion that must be executed serially. The theoretical speedup $S$ achievable when distributing a task across multiple processors is given by:
$S=\frac{1}{(1-f)+\frac{f}{s}}$ (7)
Appendix C Properties
We can now examine several useful properties of this speed-up bound $S(N,C)$. These propositions formalize the intuition that removing bottlenecks promotes generalists, never specialists, since only generalists benefit from concurrent access to subtasks.
Proposition 1 (Monotonicity). $S(N,B)$ is non-decreasing in each $C_{i}$.
Proof. $s_{i}(N,C_{i})$ is non‑decreasing in $C_{i}$ and appears in the denominator of $S$; increasing any $C_{i}$ weakly decreases the denominator and therefore weakly increases $S$.
Proposition 2 (Full concurrency favors generalists). Assume negligible task–switch costs. If $min_{i}C_{i}\geq N$ (every subtask can accommodate all agents), then $S(N,B)=N$ and any throughput‑maximizing policy set satisfies $\mathrm{SI}=0$.
Proof. When $C_{i}\geq N$ for all $i$, we have $s_{i}(N,C_{i})=N$. Substituting into the definition of $S$ gives us $S(N,B)=1/\!\sum_{i}f_{i}/N=N$. Because all agents can execute every subtask concurrently, restricting any agent’s action set cannot improve throughput; the optimum is therefore achieved by identical (generalist) policies, for which $\mathrm{SI}=0$.
Remark (Switching costs). When subtask times vary across workstations or switching between subtasks incurs large delays, the guarantee for generalist policies no longer holds; specialists may achieve higher throughput.
Proposition 3 (Bottlenecks induce specialization). If $S(N,B)<N$, then no fully generalist team can attain throughput proportional to $N$; a role allocation with $\mathrm{SI}>0$ is required for optimality.
Proof. $S<N$ implies $\exists\,i$ with $C_{i}<N$. Under identical generalist policies, the expected number of agents attempting subtask $i$ at any instant is $N$, but at most $C_{i}$ can make progress; the remainder are idle. This idle fraction reduces realized speed‑up below $N$. A specialist allocation that assigns at most $C_{i}$ agents to subtask $i$ and redistributes the rest to other subtasks eliminates that idle time. Hence any throughput‑optimal policy set must have $\mathrm{SI}>0$.
Appendix D Layout Graphs
@@
$\mathcal{V}=\{u\mid u\in\{(i,j)\mid i\in[1,n],j\in[1,m]\}\setminus\text{obstacles}\}$
@@
We represent spatial and resource bottlenecks as two global sets over $G=(\mathcal{V},\mathcal{E})$: $C^{s}\subseteq\mathcal{E}$, $C^{r}\subseteq\mathcal{V}$, where $C^{s}$ consists of edges in $\mathcal{E}$ that constrain movement and $C^{r}$ consists of nodes in $\mathcal{V}$ with task-relevant objects (e.g., workstations). Each subtask $i$ is constrained by a subset of bottlenecks relevant to the path ($C^{s}_{i}$) and resources ($C^{r}_{i}$) required to complete it.
@@
To quantify spatial bottlenecks, we compute edge betweenness centrality, a graph-theoretic measure of congestion where high-centrality edges indicate critical bottlenecks in connectivity [40]. The betweenness centrality of an edge $e$ in graph $G=(\mathcal{V},\mathcal{E})$ is:
$B(e)=\sum_{s\neq t\in V}\frac{\phi(s,t|e)}{\phi(s,t)},$ (8)
@@
where $\phi(s,t)$ is the number of shortest paths between nodes $s$ and $t$, and $\phi(s,t|e)$ is the number of those paths that pass through edge $e$. Since not all bottlenecks affect every subtask, we compute a per-subtask spatial bottleneck score by summing the centralities of all edges along the shortest path required for that subtask: $C^{s}_{i}=\sum_{e\in\text{path}(u,v)}B(e)$, where path(u,v)(u,v) is the shortest path between subtask-relevant nodes $u$ and $v$. A higher $C^{s}_{i}$ indicates stronger spatial constraints that may inhibit parallel agent movement.
@@
To quantify resource bottlenecks, we simply sum the number of task-relevant nodes that are required for a subtask $i$. For example, in a manufacturing system there may be only two welding stations; thus, only two agents may complete that subtask at the same time. Formally, for each subtask $i$, we define a resource bottleneck score $C^{r}_{i}$ as: $C^{r}_{i}=\sum_{v\in C^{r}_{i}}c(v)$, where $c(v)$ is the capacity of each task-relevant node (or the maximum agents that can use it at once).
Appendix E Training Procedures